OpenStudy (malcolmmcswain):

List out all the subsets of A = {2, 4, blue}

OpenStudy (malcolmmcswain):

@surjithayer the suspense is killing me...

OpenStudy (anonymous):

\[\left\{ \right\},\left\{ 2 \right\},\left\{ 4 \right\},\left\{ blue \right\},\left\{ 2,4 \right\},\left\{ 2,blue \right\},\left\{ 4,blue \right\},\left\{ 2,4,blue \right\}\]

OpenStudy (malcolmmcswain):

Thank you!

OpenStudy (anonymous):

yw